OpenStudy (anonymous):

simplify using common denominator: a) sec^2 theta + cosec^2theta

OpenStudy (mertsj):

\[\sec ^2\theta +\csc ^2\theta =\frac{1}{\cos ^2 \theta}+\frac{1}{\sin ^2\theta}=\frac{\sin ^2\theta +\cos ^2\theta}{\sin^2 \theta \cos^2 \theta}\]=

OpenStudy (mertsj):

\[\frac{1}{\sin ^2\theta \cos ^2\theta}\]

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

could i ask another question?

OpenStudy (mertsj):

sure